A farmed painting weighs 30kg. The two wires supporting it makes an angle of 120° with each other. What are the tensions in the wire?


1
Expert's answer
2021-02-01T12:41:27-0500

"2T\cos60°=mg\to T=\frac{mg}{2\cos60°}=\frac{30\cdot 9.8}{2\cos60°}=294\ (N)" . Answer